Assume the solubility product of Mg(OH)_2 is 1.7 t
Answer #1

Mg(OH)   \rightarrow   Mg2+ + 2OH-
   x 2x

Ksp of Mg(OH)2 = 1.7*10-11

[Mg2+] = 1.0*10-10 M

Ksp = [Mg2+][OH-]2

1.7*10-11 = 1.0*10-10 [OH-]2

[OH-]2 = 0.17

[OH-] = 0.412 M